A right circular cylinder has a surface area of 112pi. If the height of the cylinder is 10 find the radius of its base. Show your work.
Answer by richwmiller(17219)

formula for surface area A = 2 pi r^2 + 2 pi r h
with the ends included
A = 2 pi r^2 + 2 pi r h
= 2 pi r ( r + 20 )=112pi
56=r^2+20r
r^2+20r-56=0
this gets messy so I will assume we are not including the ends (bases)
or with out the ends
A = 2 pi r h=
A = 2 pi r 10=112pi
20r=112
r=112/20=56/10=5.6